# Bi-ConvLSTM: An Ultra-Lightweight Efficient Model for Human Activity Recognition on Resource Constrained Devices

### v1

@@ Abstract
Human Activity Recognition (HAR) on resource-constrained wearables requires models that balance accuracy against strict memory and computational budgets. State of the art lightweight architectures such as TinierHAR (34K parameters) and TinyHAR (55K parameters) achieve strong accuracy, but exceed memory budgets of microcontrollers with limited SRAM once operating system overhead is considered.
We present Bi-ConvLSTM, an ultra-lightweight convolutional recurrent architecture achieving 11.4K parameters on average through two-stage convolutional feature extraction with 4 temporal pooling and a single bidirectional LSTM layer. This represents 2.9 parameter reduction versus TinierHAR and 11.9 versus DeepConvLSTM while preserving linear complexity.
Evaluation across eight diverse HAR benchmarks shows that Bi-ConvLSTM maintains competitive performance within the ultra-lightweight regime: 93.41% macro F1 on UCI-HAR, 94.46% on SKODA assembly gestures, and 88.98% on Daphnet gait freeze detection. Systematic ablation reveals task dependent component contributions where bidirectionality benefits episodic event detection, but provides marginal gains on periodic locomotion. INT8 post training quantization incurs only 0.21% average F1-score degradation, yielding a 23.0 KB average deployment footprint suitable for memory constrained edge devices.
@@ I Introduction
Human Activity Recognition (HAR) from wearable inertial sensors enables critical applications across healthcare, fitness, and ubiquitous computing [9, 6]. Continuous freeze of gait detection for Parkinson’s disease patients [15], real time fall detection for elderly care [17], and always on fitness tracking [10] share a common requirement of deployment on resource constrained wearable devices where battery life and form factor are important.
@@
Quantization analysis showing 0.21% average INT8 degradation within a 23.0 KB average deployment footprint suitable for memory constrained devices.
@@ II-E Positioning of Bi-ConvLSTM
Despite these advances, a gap persists for sub-15K parameter models compatible with the most constrained microcontrollers. Table I compares architectural choices. Bi-ConvLSTM adheres to the proven convolutional recurrent paradigm while achieving 4 temporal compression through pooling, substantially reducing LSTM sequence length. Unlike attention based approaches including Transformers, strict complexity is maintained, with the recurrent inductive bias that aligns with sequential motion data. Unlike depthwise separable convolutions, standard convolutions preserve cross channel information flow critical for low dimensional sensor data (3-9 channels typical in smartphone HAR).
| Model | Conv Type | Temporal | Attention | Params | Complexity |
@@
| µBi-ConvLSTM | Standard | BiLSTM | No | 11.4K |
@@ III-A Design Principles
P2: Standard Convolutions for Low Dimensional Input. For sensor data with few channels (3-9 typical), standard convolutions provide superior cross channel feature extraction versus depthwise separable variants. Depthwise operations limit cross channel information flow, critical when accelerometer and gyroscope channels contain complementary motion cues.
@@ III-B Architecture Overview
Bi-ConvLSTM comprises five stages optimized for minimal overhead while preserving representational capacity, as illustrated in Figure 1.
@@ III-D Bidirectional LSTM
Design Rationale: BiLSTM was selected over alternatives for three reasons: (1) bidirectionality captures temporal asymmetry in episodic events where onset and recovery patterns differ; (2) LSTM’s separate cell state provides superior gradient flow versus GRU for longer compressed sequences ( = 32 timesteps); (3) last timestep aggregation avoids attention overhead while the final bidirectional state synthesizes both full forward sequence context and complete backward information.
@@ V-A Predictive Performance
| Dataset | µBi-ConvLSTM | DeepConvLSTM | TinyHAR | TinierHAR |
@@
Figure 2 visualizes performance distributions. Despite 2.9 fewer parameters than TinierHAR, Bi-ConvLSTM maintains competitive accuracy with tight variance on UCI-HAR and SKODA, both well structured datasets with clear activity boundaries.
@@ V-B Computational Efficiency
| Metric | µBi-ConvLSTM | DeepConvLSTM | TinyHAR | TinierHAR |
| Parameters | 11.4K | 135.6K | 55.1K | 33.5K |
| MACs | 485K | 15.51M | 9.29M | 1.73M |
| F1/K-Params | 7.34 | 0.63 | 1.56 | 2.61 |
| F1/M-MACs | 172.5 | 5.54 | 9.27 | 50.5 |
| Complexity |
@@
| Dataset | µBi-ConvLSTM | DeepConvLSTM | TinyHAR | TinierHAR |
@@ V-C Quantization Analysis
This quantization resilience likely stems from our architectural simplicity: batch normalization after each convolution constrains activation ranges, the single LSTM layer avoids error accumulation through stacked quantized operations, and the 4 temporal pooling reduces the number of quantized computations. Figure 5 visualizes this robustness.
V-D Memory Footprint Analysis
@@
| Dataset | µBi-ConvLSTM | TinierHAR | TinyHAR | DeepConvLSTM |
@@
Notably, Bi-ConvLSTM’s footprint scales gracefully with input dimensionality: the 79 channel Opportunity dataset requires only 32.4 KB (41% increase from the 3 channel minimum), whereas TinyHAR increases to 305 KB (411% increase) due to attention’s quadratic channel scaling. This characteristic makes Bi-ConvLSTM particularly suitable for multi-sensor wearable platforms where memory budgets are fixed, but sensor configurations may vary.
@@ VI Ablation Studies
Critical Aggregation Choice: Mean pooling (A4) causes catastrophic 52.43% degradation on UniMiB while minimally affecting WISDM (0.21%). UniMiB’s fall detection requires identifying brief transient events where the final bidirectional state captures critical endpoint information lost through averaging.
@@
Hyperparameter sensitivity analysis (varying convFilters 8-24, lstmHidden 16-32) shows the base configuration (16 filters, 24 hidden) sits on the Pareto frontier. Scaling to 24/32 yields only 0.79% average F1 improvement across datasets while nearly doubling parameters (from approximately 10.5K to 19.3K), confirming diminishing returns beyond the base configuration.
@@ VI-A Robustness Analysis
Figure 8 provides holistic comparison. Bi-ConvLSTM achieves strong performance in efficiency dimensions while accepting modest accuracy tradeoffs, demonstrating suitability for deployment scenarios where memory and computational constraints dominate.
@@ VII-B Limitations
High Variance on Challenging Datasets: WISDM results show 12.4% standard deviation, likely due to the dataset’s lower sampling rate (20 Hz) and minimal channel count (3) creating challenging learning dynamics. This suggests sensitivity to data quality in the ultra-lightweight regime.
@@ VII-C Deployment Considerations
Bi-ConvLSTM’s compact footprint (an average of 23.0 KB INT8 ) positions it for memory constrained embedded platforms. The architecture’s complexity ensures predictable and efficient resource usage across varying sequence lengths, avoiding the quadratic scaling of attention-based alternatives.
The natural deployment strategy: Bi-ConvLSTM for single sensor wearable applications requiring extreme parameter efficiency, with larger models reserved for multi-sensor configurations where capacity limitations become apparent.

### v3

@@ Abstract
Human Activity Recognition (HAR) on resource constrained wearables requires models that balance accuracy against strict memory and computational budgets. State of the art lightweight architectures such as TinierHAR (34K parameters) and TinyHAR (55K parameters) achieve strong accuracy, but exceed memory budgets of microcontrollers with limited SRAM once operating system overhead is considered.
We present Bi-ConvLSTM, an ultra-lightweight convolutional recurrent architecture achieving 11.4K parameters on average through two stage convolutional feature extraction with 4 temporal pooling, and a single bidirectional LSTM layer. This represents 2.9 parameter reduction versus TinierHAR and 11.9 versus DeepConvLSTM while preserving linear complexity.
Evaluation across eight diverse HAR benchmarks shows that Bi-ConvLSTM maintains competitive performance within the ultra-lightweight regime: 93.41% macro F1 on UCI-HAR, 94.46% on SKODA assembly gestures, and 88.98% on Daphnet gait freeze detection. Systematic ablation reveals task dependent component contributions where bidirectionality benefits episodic event detection, but provides marginal gains on periodic locomotion. On-device deployment on the Raspberry Pi Pico 2 and ESP32 validates hardware viability under both INT8 quantized and FP32 full-precision paths. Under INT8 quantization, Bi-ConvLSTM is the only architecture achieving full 8/8 dataset coverage on both platforms, with 72.8 ms average latency on Pico 2 and 97.9% PyTorch parity on ESP32. Under FP32 deployment, it achieves 100.0% parity on all successful configurations (8/8 Pico 2, 7/8 ESP32), confirming that all INT8 fidelity degradation is a quantization artifact rather than an architectural limitation.
@@ I Introduction
Human Activity Recognition (HAR) from wearable inertial sensors enables important applications across healthcare, fitness, and ubiquitous computing [9, 6]. Continuous freeze of gait detection for Parkinson’s disease patients [15], real time fall detection for elderly care [17], and always on fitness tracking [10] share a common requirement of deployment on resource constrained wearable devices where battery life and form factor are important.
@@
Cross platform on-device validation on the Raspberry Pi Pico 2 (Arm Cortex-M33) and ESP32 (Xtensa LX6) under both INT8 quantized and FP32 full precision paths, demonstrating full 8/8 INT8 dataset deployment with 72.8 ms and 168.5 ms average latency, respectively, and 100.0% FP32 parity on all successful configurations, serving as the only architecture in this study achieving complete or near complete coverage on both targets under both precision paths.
@@ II-E Positioning of Bi-ConvLSTM
Despite these advances, a gap persists for sub-15K parameter models compatible with the most constrained microcontrollers. Table I compares architectural choices. Bi-ConvLSTM adheres to the proven convolutional recurrent paradigm while achieving 4 temporal compression through pooling, substantially reducing LSTM sequence length. Unlike attention based approaches including Transformers, strict complexity is maintained, with the recurrent inductive bias that aligns with sequential motion data. Unlike depthwise separable convolutions, standard convolutions preserve cross channel information flow important for low dimensional sensor data (3-9 channels typical in smartphone HAR).
| Model | Conv Type | Temporal | Attention | Params () | Complexity () |
@@
| Bi-ConvLSTM | Standard | BiLSTM | No | 11.4K |
@@ III-A Design Principles
P2: Standard Convolutions for Low Dimensional Input. For sensor data with few channels (3-9 typical), standard convolutions provide superior cross channel feature extraction versus depthwise separable variants. Depthwise operations limit cross channel information flow, important when accelerometer and gyroscope channels contain complementary motion cues.
@@ III-B Architecture Overview
Bi-ConvLSTM comprises five stages optimized for minimal overhead while preserving representational capacity, as illustrated in Figure 1. Input sensor signals ( channels timesteps) pass through two convolutional blocks with batch normalization, ReLU activation, and 2 max pooling each, achieving 4 total temporal compression. A single bidirectional LSTM (hidden dimension 24) processes the compressed sequence, with the final timestep representation feeding the classification head. Parameter count varies with input channels and output classes.
@@ III-D Bidirectional LSTM
Design Rationale: BiLSTM was selected over alternatives for three reasons. (1) Bidirectionality captures temporal asymmetry in episodic events where onset and recovery patterns differ. (2) LSTM’s separate cell state provides superior gradient flow versus GRU for longer compressed sequences ( = 32 timesteps). (3) Last timestep aggregation avoids attention overhead while the final bidirectional state synthesizes both full forward sequence context and complete backward information.
@@ V-A Predictive Performance
| Dataset | Bi-ConvLSTM | DeepConvLSTM | TinyHAR | TinierHAR |
@@
Figure 2 visualizes performance distributions. Bi-ConvLSTM (leftmost in each group) maintains competitive variance despite 2.9 fewer parameters than TinierHAR, particularly on UCI-HAR and SKODA, both well structured datasets with clear activity boundaries.
@@ V-B Computational Efficiency
| Metric | Bi-ConvLSTM | DeepConvLSTM | TinyHAR | TinierHAR |
| Parameters () | 11.4K | 135.6K | 55.1K | 33.5K |
| MACs () | 485K | 15.51M | 9.29M | 1.73M |
| F1/K-Params () | 7.34 | 0.63 | 1.56 | 2.61 |
| F1/M-MACs () | 172.5 | 5.54 | 9.27 | 50.5 |
| Complexity () |
@@
| Dataset | Bi-ConvLSTM | DeepConvLSTM | TinyHAR | TinierHAR |
@@ V-C Desktop Profiling
This quantization resilience likely stems from our architectural simplicity: batch normalization after each convolution constrains activation ranges, the single LSTM layer avoids error accumulation through stacked quantized operations, and the 4 temporal pooling reduces the number of quantized computations. Figure 5 visualizes this robustness, where the near diagonal alignment demonstrates quantization robustness, with average degradation of only 0.21%.
@@
| Dataset | Bi-ConvLSTM | TinierHAR | TinyHAR | DeepConvLSTM |
@@
Bi-ConvLSTM’s footprint scales gracefully with input dimensionality: the 79 channel Opportunity dataset requires only 32.4 KB (41% increase from the 3 channel minimum), whereas TinyHAR increases to 305 KB (411% increase) due to attention’s quadratic channel scaling. This characteristic makes Bi-ConvLSTM particularly suitable for multi-sensor wearable platforms where memory budgets are fixed, but sensor configurations may vary.
V-D INT8 Quantized Deployment: Raspberry Pi Pico 2
To validate deployment feasibility beyond desktop profiling, INT8 quantized TensorFlow Lite Micro models are deployed, on the Raspberry Pi Pico 2 (RP2350, dual-core Arm Cortex-M33, 520 KB SRAM, 4 MB flash). All models are compiled with the standard pico-tflmicro runtime, and executed on the device with identical arena configurations. We report tensor arena usage, model binary size, inference latency, and PyTorch parity, which is the percentage of test samples where the on-device prediction agrees with the original FP32 PyTorch model’s prediction.
Table VIII presents family level deployment results. Bi-ConvLSTM is the only architecture that completes all 8/8 datasets on this platform, achieving the lowest average latency (72.8 ms), the smallest average model binary (275 KB), and the highest average parity (85.7%). TinierHAR fits in memory for all datasets but shows lower fidelity (54.2% average parity), and 3.1 higher latency. TinyHAR fails on one dataset (Opportunity) due to buffer allocation limits, and several successful runs show severe prediction drift. DeepConvLSTM is the clearest negative result: 7/8 datasets fail during tensor allocation, and the only successful run (Daphnet) degrades to 25.5% parity, which is effectively random for a binary task.
| Model | Runs () | Fails () | Latency (ms) () | Arena (B) () | Model (B) () | Parity (%) () |
| Bi-ConvLSTM | 8 | 0 | 72.8 | 109,630 | 275,088 | 85.7 |
| TinyHAR | 7 | 1 | 354.1 | 123,188 | 195,484 | 60.4 |
| TinierHAR | 8 | 0 | 229.3 | 183,974 | 312,433 | 54.2 |
| DeepConvLSTM | 1 | 7 | 768.8 | 244,084 | 865,304 | 25.5 |
Per dataset analysis reveals that Bi-ConvLSTM maintains strong fidelity across heterogeneous benchmarks. On UCI-HAR, it preserves 89.9% parity at 78.9 ms latency. On Opportunity (79 channels) it reaches 94.7% parity despite the increased arena footprint (134 KB). The highest parity is achieved on SKODA (95.9%), where the model’s convolutional stem aligns naturally with the repetitive gesture structure. Even on Daphnet, where the binary freeze/no-freeze boundary is sensitive to quantization noise, Bi-ConvLSTM sustains 72.0% parity versus DeepConvLSTM’s 25.5%, confirming that architectural compactness translates to more predictable on-device behavior.
The failure modes of the baselines are instructive. DeepConvLSTM’s dominant failure is tensor arena allocation exhaustion: it’s 1.9 MB model binaries, and correspondingly large intermediate tensors exceed the Pico 2’s contiguous SRAM budget before inference begins. TinyHAR’s Opportunity failure is a runtime buffer resize overflow (requested 493 KB against 302 KB available), indicating that attention based intermediate activations scale poorly under fixed arena constraints. These results confirm that the INT8 model size reported in desktop profiling (Table IV) is a necessary, but insufficient predictor of on-device viability. Actual deployment must also account for runtime tensor allocation, arena fragmentation, and kernel scratch memory.
V-E INT8 Quantized Deployment: ESP32
Deployment validation is also done on the ESP32-D0WD-V3 (single-core Xtensa LX6, 520 KB SRAM, 4 MB flash, no PSRAM), a widely deployed platform in commercial IoT. Models are deployed through the native ESP-IDF esp-tflite-micro pipeline with optimized arena sizing. TinierHAR required additional memory path tuning: single core mode with IRAM exposed as 8 bit accessible heap, and a 176 KB arena allocation. An initial deployment sweep revealed catastrophic parity collapses in several baseline quantized exports. Investigation traced the root cause to activation range loss in the original INT8 export recipe, rather than to hardware limitations. A corrected mixed quantized export using INT16 activations with INT8 weights was therefore evaluated for the affected baseline rows, and promoted where it materially improved on-device fidelity.
Table IX summarizes the corrected results. Bi-ConvLSTM again achieves full 8/8 dataset coverage, the highest average PyTorch parity (97.9%), and moderate latency (168.5 ms). On-device parity on ESP32 is considerably higher than on Pico 2 (97.9% vs. 85.7%), due to the ESP32’s wider ALU datapath, and more favorable fixed point rounding behavior in the Xtensa esp-nn kernels. TinierHAR deploys on 6/8 datasets with an improved average parity of 90.8% after the corrected quantized export, making it a credible secondary result on this platform. TinyHAR deploys on 4/8 datasets with 88.6% average parity, a notable recovery from the earlier near-zero fidelity figures, though one previously successful dataset (UCI-HAR) became memory-limited after the corrected export because the more faithful quantized graph required a larger internal SRAM allocation. DeepConvLSTM fails on all 8/8 datasets due to tensor arena allocation failure on the no-PSRAM ESP32, confirming that this architecture class is fundamentally incompatible with sub-520 KB SRAM platforms.
| Model | Runs () | Fails () | Latency (ms) () | Model (B) () | Parity (%) () |
| Bi-ConvLSTM | 8 | 0 | 168.5 | 275,088 | 97.9 |
| TinyHAR | 4 | 4 | 487.0 | 218,202 | 88.6 |
| TinierHAR | 6 | 2 | 204.5 | 312,433 | 90.8 |
| DeepConvLSTM | 0 | 8 | — | 1,758,060 | — |
Per dataset ESP32 results reveal additional deployment insights. Bi-ConvLSTM achieves near perfect parity on several benchmarks: 99.3% on UCI-HAR, 99.6% on SKODA, and 99.2% on Opportunity. These figures exceed the desktop INT8 quantization degradation bounds (Table IV), confirming that the quantized model generalizes to heterogeneous hardware backends. The Daphnet configuration achieves the lowest latency (95.2 ms) due to its smaller 144 KB arena, and reduced model binary (155 KB), displaying the favorable latency memory scaling of the architecture.
The corrected baseline results show an important methodological concern. The earlier catastrophic parity collapses observed for TinyHAR and TinierHAR were not purely hardware limitations, but were caused by activation range loss in the standard full INT8 export path. Replacing the export recipe with INT16 activations and INT8 weights recovered meaningful fidelity for TinyHAR on MotionSense (from near-zero to 88.1%), and for TinierHAR on MotionSense (to 91.6%). This recovery came at a cost: the more faithful quantized graphs sometimes required larger tensor arenas, pushing previously deployable configurations past the internal SRAM limit. TinyHAR on UCI-HAR is an example, where the corrected export fits on desktop, but fails to reserve a contiguous 256 KB arena on the ESP32.
The ESP32 failure taxonomy remains practically informative. Two distinct failure classes emerge: (1) tensor arena allocation failures, where the board cannot reserve the requested contiguous internal SRAM block at startup (dominant for DeepConvLSTM, and four TinyHAR datasets), and (2) AllocateTensors resize failures, where the arena exists but planner scratch memory exceeds the residual budget (remaining TinierHAR failures on PAMAP2 and Opportunity). This distinction matters because the second class indicates the native inference path is valid, and close to working, as it is no longer an ABI or kernel compatibility problem, but a quantitative arena budget shortfall. Figure 6 visualizes the cross platform deployment landscape, showing how the Bi-ConvLSTM occupies the ideal top left region of the Pareto space on both platforms, combining the lowest latency with the highest fidelity.
Across both platforms under INT8 quantization, Bi-ConvLSTM uniquely combines full dataset coverage with low latency, and high inference fidelity. Its 4 temporal pooling, and single layer BiLSTM, yield smaller tensor arenas, shorter inference paths, and predictable runtime under fixed memory constraints, empirically validating the efficiency claims of Sections V-A and V-B.
V-F FP32 Full-Precision Deployment
To isolate quantization induced fidelity loss from architectural deployment limitations, FP32 full-precision TFLite models are deployed, on both platforms. The FP32 path eliminates the INT8 rounding error entirely but produces larger model binaries and tensor arenas, reducing dataset coverage on memory constrained MCUs. Tables X and XI summarize the FP32 results.
| Model | Runs () | Fails () | Latency (ms) () | Model (B) () | Parity (%) () |
| Bi-ConvLSTM | 8 | 0 | 83.0 | 555,221 | 100.0 |
| TinierHAR | 7 | 1 | 231.5 | 647,156 | 100.0 |
| TinyHAR | 5 | 3 | 486.1 | 626,212 | 89.6 |
| DeepConvLSTM | 1 | 7 | 2034.6 | 1,273,108 | 26.3 |
| Model | Runs () | Fails () | Latency (ms) () | Model (B) () | Parity (%) () |
| Bi-ConvLSTM | 7 | 1 | 67.3 | 498,528 | 100.0 |
| TinierHAR | 3 | 5 | 96.4 | 514,784 | 100.0 |
| TinyHAR | 3 | 5 | 352.9 | 517,792 | 88.1 |
| DeepConvLSTM | 0 | 8 | — | — | — |
On Pico 2, Bi-ConvLSTM again achieves full 8/8 FP32 coverage with 100.0% PyTorch parity and 83.0 ms average latency, confirming that all INT8 parity degradation observed in Table VIII is due to quantization rather than to architectural or runtime limitations. TinierHAR deploys on 7/8 datasets (failing only on Opportunity due to buffer resize overflow) with perfect parity on all successful runs, demonstrating that the 54.2% INT8 parity reported in Section V-C was entirely a quantization artifact. TinyHAR achieves 5/8 FP32 coverage (failing on PAMAP2, Opportunity, and SKODA) at 89.6% average parity, where the remaining parity gap on Daphnet (64.9%) persists even in FP32, and may reflect a training convergence issue rather than a deployment artifact.
On ESP32, Bi-ConvLSTM achieves 7/8 FP32 coverage (failing only on Opportunity, where the unquantized arena exceeds 185 KB internal SRAM) with 100.0% parity on all successful datasets, and 67.3 ms average latency, which is 2.5 faster than the INT8 path on the same platform. This latency reduction occurs because the ESP32’s Xtensa core natively supports single-precision floating point operations through hardware, making FP32 inference faster than the software emulated INT8 dequantize-compute path in esp-tflite-micro. TinierHAR and TinyHAR each deploy on only 3/8 FP32 datasets (WISDM, UniMiB, and Daphnet), with TinierHAR achieving perfect parity on its successful configurations. DeepConvLSTM fails on all 8/8 datasets in FP32 as well, confirming that its memory footprint is fundamentally incompatible with both platforms regardless of precision.
The FP32 results provide three key insights. First, the INT8 parity degradation observed in Sections V-C and V-D is confirmed to be a quantization artifact rather than an architectural deployment limitation, because all models that deploy in FP32 achieve higher fidelity. Second, FP32 coverage is lower than INT8 coverage for every architecture except Bi-ConvLSTM on Pico 2 (where both achieve 8/8), showing the fundamental coverage fidelity tradeoff in TinyML deployment. Third, Bi-ConvLSTM is the only architecture achieving near complete FP32 coverage on both platforms (8/8 Pico 2, 7/8 ESP32), confirming that its compact tensor arena footprint provides deployment headroom that larger architectures lack even when quantization overhead is removed. Figure 7 visualizes the FP32 deployment landscape.
@@ VI Ablation Studies
Important Aggregation Choice: Mean pooling (A4) causes catastrophic 52.43% degradation on UniMiB while minimally affecting WISDM (0.21%). UniMiB’s fall detection requires identifying brief transient events where the final bidirectional state captures important endpoint information lost through averaging.
@@
As shown in Figure 9, the base configuration achieves best or near best performance on 5 of 8 datasets. Hyperparameter sensitivity analysis (varying convFilters 8-24, lstmHidden 16-32) shows the base configuration (16 filters, 24 hidden) sits on the Pareto frontier. Scaling to 24/32 yields only 0.79% average F1 improvement across datasets while nearly doubling parameters (from approximately 10.5K to 19.3K), confirming diminishing returns beyond the base configuration.
@@ VI-A Robustness Analysis
Figure 10 provides a comparison. All metrics are scaled to [0,1] where larger values indicate better performance. Bi-ConvLSTM achieves the largest area in efficiency dimensions while accepting modest accuracy tradeoffs, demonstrating suitability for deployment scenarios where memory and computational constraints dominate.
@@ VII-B Limitations
High Variance on Challenging Datasets: WISDM results show that 12.4% standard deviation, likely due to the dataset’s lower sampling rate (20 Hz), and minimal channel count (3) creating challenging learning dynamics. This suggests sensitivity to data quality in the ultra-lightweight regime.
@@ VII-C Deployment Considerations
The on-device results across both INT8 and FP32 paths (Sections V-C through V-E) change the deployment discussion from theoretical projection to empirical evidence. Under INT8 quantization, Bi-ConvLSTM achieves full 8/8 dataset coverage on both the Pico 2 (72.8 ms average, with 85.7% parity) and ESP32 (168.5 ms average, with 97.9% parity). Under FP32 full precision deployment, Bi-ConvLSTM achieves 8/8 on Pico 2 and 7/8 on ESP32, both with 100.0% PyTorch parity, confirming that all INT8 fidelity degradation is a quantization artifact. No other architecture achieves full coverage on both platforms under either precision path. DeepConvLSTM is non-deployable on both platforms in both INT8 and FP32 despite strong desktop accuracy.
The FP32 results isolate quantization as the sole source of the parity gaps reported in Sections V-C and V-D. TinierHAR’s INT8 parity on Pico 2 (54.2%) recovers to 100.0% in FP32, confirming that the CMSIS-NN INT8 kernel path is the bottleneck rather than any architectural deployment limitation. The corrected ESP32 INT8 baselines add further nuance: replacing the full INT8 export recipe with INT16 activations and INT8 weights recovered considerable fidelity for TinierHAR (90.8%) and TinyHAR (88.6%), though the corrected exports sometimes traded fidelity for increased memory, pushing previously deployable configurations past the internal SRAM limit.
The INT8 parity gap between platforms (85.7% on Pico 2 vs. 97.9% on ESP32) highlights that on-device quantized inference fidelity is backend dependent: the ESP32’s Xtensa esp-nn kernels produce more favorable fixed point rounding than the Cortex-M33’s CMSIS-NN path. This cross platform variation is an underreported concern in TinyML research, where desktop INT8 evaluation is often treated as a reliable proxy for on-device behavior. The FP32 deployment path eliminates this variation entirely, but at the cost of reduced coverage due to larger tensor arenas.
The natural deployment strategy depends on the application constraint. For maximum coverage with acceptable fidelity, INT8 deployment of Bi-ConvLSTM provides guaranteed 8/8 coverage on both platforms. For maximum fidelity on platforms with sufficient memory, FP32 deployment of Bi-ConvLSTM achieves perfect parity on all deployable configurations. For applications where partial coverage is acceptable, TinierHAR in FP32 represents a viable alternative on platforms with sufficient IRAM budget.